# AWARE: Audio Watermarking via Adversarial Resistance to Edits

### v1

@@ 6 Ablation Studies
To evaluate effectiveness and rationale of key architectural and representational design choices, we conducted a series of ablation studies. Each experiment isolates one architectural or representational component while keeping all other settings fixed. This allows us to examine how each component, from the Bitwise Readout Head to kernel size and spectral domain selection, impacts robustness and consistency under typical audio distortions. For every ablation, we chose representative attack subsets that the ablated component is expected to affect. The results are summarized in the tables below as mean BER across those cases.
@@
In Table 3, we compare the proposed BRH with a variant that replaces it by a conventional FC output layer, which binds detection to absolute positions and thus loses time-order invariance. We focus on sample deletion and time-stretch conditions, as they most directly probe temporal robustness.
Table 3: BER (%) under deletion (SD) and time-stretch (TS) attacks for models with and without BRH.
Model SD TS
w/ BRH (proposed) 3.74 5.53
w/o BRH (FC layer) 30.91 7.84
The BRH substantially improves robustness under temporal desynchronization, as it aggregates evidence over time without depending on frame order or absolute position. This effect is particularly evident in fragment-level detection, where portions of the signal are entirely removed: the BRH can still accumulate sufficient evidence from the remaining segments to correctly infer watermark bits, while the FC variant collapses once the input continuity is broken.
6.2 Kernel Size 1 vs. 3
Next, we assess the effect of convolutional kernel size along the temporal axis. We analyze convolutional kernels of size $1$ and $3$. Kernel size $1$ avoids temporal mixing and maintains activation stability under cuts, whereas kernel size $3$ introduces context leakage across frames.
Table 4: BER (%) under temporal edits for different convolutional kernel sizes.
Kernel size SD TS
11 (proposed) 3.74 5.53
33 6.06 13.96
Results, given in Table 4, support the choice of minimal temporal context in feature extraction blocks. Restricting convolutional kernels to size $1$ keeps activations independent across frames and preserves stability under deletions and temporal warping.
@@
Finally, we compare the proposed Mel-based detector with a variant operating directly on STFT magnitudes. The Mel front-end performs perceptual band aggregation, which stabilizes frequency-domain statistics and aligns the representation with human-critical bands and common synthesis pipelines. Accordingly, we probe attacks that directly stress these properties: LPF/HPF, MP3, and neural vocoder resynthesis, precisely where Mel aggregation is expected to enhance the robustness.
Table 5: BER (%) under neural vocoder (NV) resynthesis and spectral/compression attacks with and without the Mel front-end.
Front-End NV LPF HPF MP3
w/ Mel (proposed) 1.61 0.00 0.0 0.71
w/o Mel (STFT only) 50.30 0.68 0.83 1.42
Results in Table 5 suggest that for spectral edits (LPF/HPF) and compression (MP3), absolute differences are modest, yet consistently favor the Mel front-end. Under NV resynthesis, the gap becomes substantial. The Mel-based detector remains highly reliable, whereas the STFT-only variant degrades sharply.
@@ 7 Conclusion
This paper advocates robustness by design for digital audio watermarking and introduces a time–order–agnostic detector with a Bitwise Readout Head. Adversarial evaluation serves as a diagnostic tool to surface right inductive biases, without relying on heavy attack simulation. The system delivers high audio quality and consistently low BER across diverse edits, often surpassing strong learning-based baselines. However, rather than pursuing “state-of-the-art” claims, particularly tenuous in this domain, given competing metrics and trade-offs, we aim to provide principled guidance on design choices tailored to the intended threat model for digital audio watermarking use cases.

### v2

@@ 7 Ablation Studies
To evaluate effectiveness and rationale of key architectural and representational design choices, we conduct ablation studies. Each experiment isolates one architectural or representational component while keeping all other settings fixed. This allows us to examine how each component impacts robustness and consistency under typical audio distortions. For every ablation, we chose representative attack subsets that the ablated component is expected to affect. The results are summarized in the tables below as mean BER across those cases.
@@
Table 3 ablates the BRH, as a key architetural component, against a conventional FC output layer, preceded by global pooling over the temporal dimension. BRH markedly improves performance on edits that disrupt alignment or ordering (TS, PA, RA and VT), as it aggregates evidence over time without relying on absolute positions. By contrast, the FC variant collapses once the input continuity is broken.
Table 3: BER with and without BRH.
Model PS NAC TS SR PA RA VT
w/ BRH 0.008 0.179 0.016 0.095 0.000 0.022 0.095
w/o BRH 0.359 0.470 0.373 0.170 0.386 0.404 0.170
@@
Finally, we compare the proposed Mel-based detector with a variant operating directly on STFT magnitudes. The Mel front-end performs perceptual band aggregation, which stabilizes frequency-domain statistics and aligns the representation with human-critical bands and common synthesis pipelines. Accordingly, we probe attacks that directly stress these properties: LPF/HPF, MP3, NV, RA and VP, precisely where Mel aggregation is expected to enhance the robustness.
Table 4: BER with and without the Mel front-end.
Front-End LPF HPF MP3 PS NV RA VT
w/ Mel (proposed) 0.000 0.000 0.001 0.014 0.016 0.024 0.102
w/o Mel (STFT only) 0.007 0.008 0.014 0.147 0.530 0.058 0.136
Results in Table 4 suggest that for spectral edits (LPF/HPF) and compression (MP3), absolute differences are modest, yet consistently favor the Mel front-end. Under PS and NV resynthesis, the gap becomes substantial. The Mel-based detector remains highly reliable, whereas the STFT-only variant degrades sharply. A clear, though more moderate, advantage is also observed under RA, which indicates that Mel aggregation provides additional robustness to acoustic channel coloration and recording/playback artifacts. Under VT, the Mel front-end again improves performance, and shows better tolerance to the compound distortions introduced by networked speech communication pipelines.
@@ 8 Conclusion
This paper advocates robustness by design for digital audio watermarking and introduces a time–order–agnostic detector with a Bitwise Readout Head. The proposed system achieves high audio quality and low error rate across a wide range of audio edits, often surpassing strong learning-based baselines despite not being trained on attack-specific distortions. However, rather than pursuing “state-of-the-art” claims, particularly tenuous in this domain, given competing metrics and trade-offs, we aim to provide principled guidance on design choices aligned with practical threat models for digital audio watermarking use cases.